# Synesthesia of Machines (SoM)-Enhanced Wideband Multi-User CSI Learning With LiDAR Sensing

### v2

@@ IV Simulation Results
In this section, we present simulation results to evaluate the performance of the proposed LE-CLN. In our simulations, both the wireless channel data and LiDAR data are derived from the vehicular urban crossroad scenario in the M3SC dataset [18], the BS is equipped with $N_{\rm{t}}=32$ antennas. The carrier frequency is $f_{\rm{c}}=28$GHz and the bandwidth is $100$MHz consisting of $N_{\rm{s}}=64$ subcarriers. We set the number of multiple paths as $n_{\rm{p}}=8$ and the heights of transceivers as $h_{\rm{t}}=5$m and $h_{\rm{r}}=1$m. The size of the over-sampled subspace $N_{\rm{w}}$ is set to $13$. The parameters of LiDAR are: $a_{\rm{h}}=0.36^{\circ}$, $f_{\rm{up}}=15^{{\circ}}$, $f_{\rm{down}}=-25^{{\circ}}$, $M_{\rm{b}}=64$. The resolution of the ULO-DFT codebook is set to $D=2N_{\rm{t}}=64$. The quantization bit number of the phase shifters adopted at BS is set to $B=3$, resulting in the quantization of the RF phase into $Q\triangleq 2^{B}$ discrete values. Each element of $\mathbf{F}_{\rm{R}}$ is randomly selected from the set $\{{e^{\frac{j2\pi n}{Q}},n=1,2,\cdots,Q}\}$. Table I lists the hyper-parameters used for fine-tuning the LE-CLN model. All simulation results presented are averaged over $300$ independent realizations.
@@
The proposed LE-CLN scheme is compared with the existing OMP [19], AMP [20], LS [21], CENN [8], and GM-LAMP [10] schemes. The normalized mean square error (NMSE) is used as the estimation performance metric given by ${\rm NMSE}={{\|\hat{\mathbf{H}}-\mathbf{H}\|}_{\rm{F}}^{2}}/{{\|\mathbf{H}\|}_{\rm{F}}^{2}}$. The number of measurements consumed by each scheme in the time domain is indicated in the legend. In the frequency domain, $8$ pilots are used and uniformly distributed across all subcarriers in all schemes. The average received SNR is defined as $\frac{\|\mathbf{F}_{\rm{R}}\mathbf{H}_{k}\|_{\rm{F}}^{2}}{\|\mathbf{F}_{\rm{R}}\|_{\rm{F}}^{2}\mathbb{E}\{\mathbf{N}_{k}^{2}\}}$.
@@
In Fig. 3c, we compare the upper bound of the downlink spectral efficiency (SE) of different schemes, which is measured by the fully-digital precoding. The zero-forcing (ZF) precoder at the $m$-th subcarrier can be calculated by $\mathbf{f}_{k}^{m}=[(\hat{\mathbf{h}}_{k}^{m})^{H}\hat{\mathbf{h}}_{k}^{m}]^{-1}(\hat{\mathbf{h}}_{k}^{m})^{H}$. The power of $\mathbf{f}_{k}^{m}$ is normalized by $\|\mathbf{f}_{k}^{m}\|_{\rm{2}}=1$. According to the specifications on the fifth-generation New Radio sidelink [22] and the channel coherence time in vehicular networks, we assume that total number of symbols is $900$, consisting of $N_{\rm{P}}$ pilot symbols and $N_{\rm{D}}$ data symbols. Then, the SE is given by $R_{k}=\frac{N_{\rm{D}}}{900}\frac{1}{N_{\rm{s}}}\sum_{m=1}^{N_{\rm{s}}}\log_{2}{(1+\frac{|\mathbf{f}_{k}^{m}\mathbf{h}_{k}^{m}|^{2}}{\sigma^{2}})}$ in the considered OFDM system. To clearly demonstrate the SE performance differences, Fig. 3c shows the normalized SE with the proposed LE-CLN(8)’s SE being the basis. The normalized SE is defined by the ratio between SE achieved by all schemes and that achieved by the proposed LE-CLN(8) scheme. A ratio greater than $1$ indicates that the SE achieved by a certain scheme is higher than that achieved by LE-CLN(8). As can be seen, LE-CLN(8) achieves the highest SE in high SNRs ($[12,21]$dB). At low SNRs ($[-3,12]$dB), the LE-CLN(16), LE-CLN(32), LS(32), and CENN(32) schemes achieve higher SE performance owing to more accurate channel estimation results. Among these, the LE-CLN(32) scheme demonstrates the most optimal performance.
@@ V Conclusion
In this study, we investigated the use of LiDAR to improve channel estimation accuracy in wideband multi-user MIMO-OFDM systems. The LE-CLN model extracts additional angular features from pilots using the ULO-DFT codebook, which is constructed based on user positioning information. Furthermore, by integrating out-of-band wireless channel characteristics obtained from LiDAR environmental detection, LE-CLN achieves comparable or even superior estimation accuracy compared to standard benchmarks while reducing the number of required pilots. This performance enhancement is particularly pronounced in scenarios with lower SNRs and limited pilot measurements. Additionally, simulation results demonstrate that LE-CLN can dynamically adjust LiDAR utilization based on varying channel conditions, further enhancing its overall performance.

### v3

@@ IV Simulation Results
In this section, we present simulation results to evaluate the performance of the proposed LE-CLN. In our simulations, both the wireless channel data and LiDAR data are derived from the vehicular urban crossroad scenario in the M3SC dataset [18]. The BS is equipped with $N_{\rm{t}}=32$ antennas. The carrier frequency is $f_{\rm{c}}=28$GHz and the bandwidth is $100$MHz consisting of $N_{\rm{s}}=64$ subcarriers. We set the number of multiple paths as $n_{\rm{p}}=8$ and the heights of transceivers as $h_{\rm{t}}=5$m and $h_{\rm{r}}=1$m. The size of the over-sampled subspace $N_{\rm{w}}$ is set to $13$. The parameters of LiDAR are: $a_{\rm{h}}=0.36^{\circ}$, $f_{\rm{up}}=15^{{\circ}}$, $f_{\rm{down}}=-25^{{\circ}}$, $M_{\rm{b}}=64$. The resolution of the ULO-DFT codebook is set to $D=2N_{\rm{t}}=64$. The quantization bit number of the phase shifters adopted at BS is set to $B=3$, resulting in the quantization of the RF phase into $Q\triangleq 2^{B}$ discrete values. Each element of $\mathbf{F}_{\rm{R}}$ is randomly selected from the set $\{{e^{\frac{j2\pi n}{Q}},n=1,2,\cdots,Q}\}$. Table I lists the hyper-parameters used for fine-tuning the LE-CLN model. All simulation results presented are averaged over $300$ independent realizations.
@@
The proposed LE-CLN scheme is compared with the existing OMP [19], AMP [20], LS [21], CENN [8], and GM-LAMP [10] schemes. The normalized mean square error (NMSE) is used as the estimation performance metric given by ${\rm NMSE}={{\|\hat{\mathbf{H}}-\mathbf{H}\|}_{\rm{F}}^{2}}/{{\|\mathbf{H}\|}_{\rm{F}}^{2}}$. The number of measurements consumed by each scheme in the time domain is indicated in the legend. In the frequency domain, $8$ pilots are used and uniformly distributed across all subcarriers in all schemes. The average received SNR is defined as ${\rm SNR}={\|\mathbf{F}_{\rm{R}}\mathbf{H}_{k}\|_{\rm{F}}^{2}}/({\sigma^{2}\|\mathbf{F}_{\rm{R}}\|_{\rm{F}}^{2}})$.
@@
In Fig. 3c, we compare the upper bound of the downlink spectral efficiency (SE) of different schemes, which is measured by the fully-digital precoding. The zero-forcing (ZF) precoder at the $m$-th subcarrier can be calculated by $\mathbf{f}_{k}^{m}=[(\hat{\mathbf{h}}_{k}^{m})^{\rm H}\hat{\mathbf{h}}_{k}^{m}]^{-1}(\hat{\mathbf{h}}_{k}^{m})^{\rm H}$. The power of $\mathbf{f}_{k}^{m}$ is normalized by $\|\mathbf{f}_{k}^{m}\|_{\rm{2}}=1$. According to the specifications on the fifth-generation New Radio sidelink [22] and the channel coherence time in vehicular networks, we assume that total number of symbols is $900$, consisting of $N_{\rm{P}}$ pilot symbols and $N_{\rm{D}}$ data symbols. Then, the SE is given by $R_{k}=\frac{N_{\rm{D}}}{900}\frac{1}{N_{\rm{s}}}\sum_{m=1}^{N_{\rm{s}}}\log_{2}{(1+{|\mathbf{f}_{k}^{m}\mathbf{h}_{k}^{m}|^{2}}/{\sigma^{2}})}$ in the considered OFDM system. To clearly demonstrate the SE performance differences, Fig. 3c shows the normalized SE with the proposed LE-CLN(8)’s SE being the basis. The normalized SE is defined by the ratio between SE achieved by all schemes and that achieved by the proposed LE-CLN(8) scheme. A ratio greater than $1$ indicates that the SE achieved by a certain scheme is higher than that achieved by LE-CLN(8). As can be seen, LE-CLN(8) achieves the highest SE in high SNRs ($[12,21]$dB). At low SNRs ($[-3,12]$dB), the LE-CLN(16), LE-CLN(32), LS(32), and CENN(32) schemes achieve higher SE performance owing to more accurate channel estimation results. Among these, the LE-CLN(32) scheme demonstrates the most optimal performance.
@@ V Conclusion
In this study, we investigated the use of LiDAR to improve channel estimation accuracy in wideband multi-user MIMO-OFDM systems. The LE-CLN model extracts additional angular features from pilots using the ULO-DFT codebook, which is constructed based on user positioning information. Furthermore, by integrating OOB wireless channel characteristics obtained from LiDAR environmental detection, LE-CLN achieves comparable or even superior estimation accuracy compared to standard benchmarks while reducing the number of required pilots. This performance enhancement is particularly pronounced in scenarios with lower SNRs and limited pilot measurements. Additionally, simulation results demonstrate that LE-CLN can dynamically adjust LiDAR utilization based on varying channel conditions, further enhancing its overall performance.